# M2Distill: Multi-Modal Distillation for Lifelong Imitation Learning

### v2

@@ Abstract
Lifelong imitation learning for manipulation tasks poses significant challenges due to distribution shifts that occur in incremental learning steps. Existing methods often focus on unsupervised skill discovery to construct an ever-growing skill library or distillation from multiple policies, which can lead to scalability issues as diverse manipulation tasks are continually introduced and may fail to ensure a consistent latent space throughout the learning process, leading to catastrophic forgetting of previously learned skills. In this paper, we introduce M2Distill, a multi-modal distillation-based method for lifelong imitation learning focusing on preserving consistent latent space across vision, language, and action distributions throughout the learning process. By regulating the shifts in latent representations across different modalities from previous to current steps, and reducing discrepancies in Gaussian Mixture Model (GMM) policies between consecutive learning steps, we ensure that the learned policy retains its ability to perform previously learned tasks while seamlessly integrating new skills. Extensive evaluations on the LIBERO lifelong imitation learning benchmark suites, including LIBERO-OBJECT, LIBERO-GOAL, and LIBERO-SPATIAL, demonstrate that our method consistently outperforms prior state-of-the-art methods across all evaluated metrics.
@@ I Introduction
Figure 1: t-SNE visualization of latent space deformation for AgentView images using Experience Replay (ER) method across two consecutive steps (i.e., $T^{2}$ and $T^{3}$) in a lifelong imitation learning scenario on LIBERO-OBJECT. The t-SNE plot highlights significant shifts in latent representations, contributing to catastrophic forgetting, while the accompanying bar plot shows reduced representation drift in M2Distill during the sequential learning of manipulation tasks.
@@ II Related Work
Catastrophic forgetting is well-studied for various problems in computer vision, including classification [17], detection [18], and semantic segmentation [19], within lifelong learning scenarios. In the literature of lifelong learning, dynamic architecture [20, 21], regularization [22, 23] and memory-replay [24, 25, 11, 6] based approaches have been proposed to tackle the catastrophic forgetting. Regularization based methods control the changes in the network’s weight by introducing new regularization terms. Memory-replay based strategies store a subset of past examples and replay with new examples.
@@ III Problem Formulation
One of the key challenges in lifelong imitation learning is that, as the robot progresses to new tasks, it loses direct access to the previous tasks $T^{1},\dots,T^{k-1}$ [29]. This limitation necessitates that the robot not only performs well on the current task but also effectively retains and transfers knowledge from prior tasks, ensuring that the policy evolves in a way that supports future learning. The goal is to balance leveraging previously acquired knowledge and adapting to new challenges, thereby optimizing the robot’s overall learning efficiency and performance over its entire operational lifespan.
@@ V Experimental Settings
TABLE I: Experimental results across three different LIBERO task suites. The reported values are averages from three seeds, including the mean and standard error. Best values are highlighted in bold, and second-best values are underlined. The dash (-) indicates failure to reproduce results.
@@
Sequential .62 ($\pm$ .00) .63 ($\pm$ .02) .30 ($\pm$ .00) .55 ($\pm$ .00) .70 ($\pm$ .01) .23 ($\pm$ .00) .72 ($\pm$ .01) .81 ($\pm$ .01) .20 ($\pm$ .01)
EWC [11] .56 ($\pm$ .03) .69 ($\pm$ .02) .16 ($\pm$ .02) .32 ($\pm$ .02) .48 ($\pm$ .03) .06 ($\pm$ .00) .23 ($\pm$ .01) .33 ($\pm$ .01) .06 ($\pm$ .01)
ER [12] .56 ($\pm$ .01) .24 ($\pm$ .00) .49 ($\pm$ .01) .53 ($\pm$ .01) .36 ($\pm$ .01) .47 ($\pm$ .02) .65 ($\pm$ .03) .27 ($\pm$ .03) .56 ($\pm$ .01)
BUDS [7] .52 ($\pm$ .02) .21 ($\pm$ .01) .47 ($\pm$ .01) .50 ($\pm$ .01) .39 ($\pm$ .01) .42 ($\pm$ .01) - - -
LOTUS [9] .74 ($\pm$ .03) .11 ($\pm$ .01) .65 ($\pm$ .03) .61 ($\pm$ .03) .30 ($\pm$ .01) .56 ($\pm$ .01) - - -
Ours .75 ($\pm$ .03) .08 ($\pm$ .05) .69 ($\pm$ .04) .71 ($\pm$ .01) .20 ($\pm$ .03) .57 ($\pm$ .02) .74 ($\pm$ .01) .11 ($\pm$ .01) .61 ($\pm$ .02)
@@ V-A Training and Implementation Details
  • ER [12]: A rehearsal-based method that preserves a memory buffer containing samples from previous tasks and uses this buffer to facilitate learning of new tasks. We impose a capacity limit of 1000 trajectories on the replay buffer.

### v3

@@ Abstract
Lifelong imitation learning for manipulation tasks poses significant challenges due to distribution shifts that occur in incremental learning steps. Existing methods often rely on unsupervised skill discovery to construct an ever-growing skill library or distillation from multiple policies, which can lead to scalability issues as diverse manipulation tasks are continually introduced and may fail to ensure a consistent latent space throughout the learning process, leading to catastrophic forgetting of previously learned skills. In this paper, we introduce M2Distill, a multi-modal distillation-based method for lifelong imitation learning focusing on preserving consistent latent space across vision, language, and action distributions throughout the learning process. By regulating the shifts in latent representations across different modalities from previous to current steps, and reducing discrepancies in Gaussian Mixture Model (GMM) policies between consecutive learning steps, we ensure that the learned policy retains its ability to perform previously learned tasks while seamlessly integrating new skills. Evaluations on the LIBERO lifelong imitation learning benchmark suites, including LIBERO-OBJECT, LIBERO-GOAL, and LIBERO-SPATIAL, demonstrate that our method consistently outperforms prior state-of-the-art methods across all evaluated metrics.
@@ I Introduction
Figure 1: t-SNE visualization of latent space deformation for AgentView images using Experience Replay (ER) method across two consecutive steps (i.e., $T^{2}$ and $T^{3}$) in a lifelong imitation learning scenario on LIBERO-OBJECT. The t-SNE plots highlight significant shifts in latent representations, contributing to catastrophic forgetting, while the accompanying bar plot shows reduced representation drift in M2Distill during the sequential learning of manipulation tasks.
@@ II Related Work
Catastrophic forgetting is well-studied for various problems in computer vision, including classification [17], detection [18], and semantic segmentation [19], within lifelong learning scenarios. In the literature of lifelong learning, dynamic architecture [20, 21], regularization [22, 23], and memory-replay [24, 25, 11, 6] based approaches have been proposed to tackle the catastrophic forgetting. Regularization based methods control the changes in the network’s weight by introducing new regularization terms. Memory-replay based strategies store a subset of past examples and replay with new examples.
@@ III Problem Formulation
One of the challenges in lifelong imitation learning is that, as the robot progresses to new tasks, it loses direct access to the previous tasks $T^{1},\dots,T^{k-1}$ [29]. This limitation necessitates that the robot not only performs well on the current task but also effectively retains and transfers knowledge from prior tasks, ensuring that the policy evolves to support future learning. The goal is to balance leveraging previously acquired knowledge and adapting to new challenges, thereby optimizing the robot’s overall learning efficiency and performance over its entire operational lifespan.
@@ V Experimental Settings
TABLE I: Experimental results across three different LIBERO task suites. The reported values are averages from three seeds, including the mean and standard error. The best values are highlighted in bold, and the second-best values are underlined. The dash (-) indicates a failure to reproduce results. All metrics are measured based on success rates (%).
@@
Sequential 62.0 ($\pm$ 1.0) 63.0 ($\pm$ 2.0) 30.0 ($\pm$ 1.0) 55.0 ($\pm$ 1.0) 70.0 ($\pm$ 1.0) 23.0 ($\pm$ 1.0) 72.0 ($\pm$ 1.0) 81.0 ($\pm$ 1.0) 20.0 ($\pm$ 1.0)
EWC [11] 56.0 ($\pm$ 3.0) 69.0 ($\pm$ 2.0) 16.0 ($\pm$ 2.0) 32.0 ($\pm$ 2.0) 48.0 ($\pm$ 3.0) 6.0 ($\pm$ 1.0) 23.0 ($\pm$ 1.0) 33.0 ($\pm$ 1.0) 6.0 ($\pm$ 1.0)
ER [12] 56.0 ($\pm$ 1.0) 24.0 ($\pm$ 1.0) 49.0 ($\pm$ 1.0) 53.0 ($\pm$ 1.0) 36.0 ($\pm$ 1.0) 47.0 ($\pm$ 2.0) 65.0 ($\pm$ 3.0) 27.0 ($\pm$ 3.0) 56.0 ($\pm$ 1.0)
BUDS [7] 52.0 ($\pm$ 2.0) 21.0 ($\pm$ 1.0) 47.0 ($\pm$ 1.0) 50.0 ($\pm$ 1.0) 39.0 ($\pm$ 1.0) 42.0 ($\pm$ 1.0) - - -
LOTUS [9] 74.0 ($\pm$ 3.0) 11.0 ($\pm$ 1.0) 65.0 ($\pm$ 3.0) 61.0 ($\pm$ 3.0) 30.0 ($\pm$ 1.0) 56.0 ($\pm$ 1.0) - - -
Ours 75.0 ($\pm$ 3.0) 8.0 ($\pm$ 5.0) 69.0 ($\pm$ 4.0) 71.0 ($\pm$ 1.0) 20.0 ($\pm$ 3.0) 57.0 ($\pm$ 2.0) 74.0 ($\pm$ 1.0) 11.0 ($\pm$ 1.0) 61.0 ($\pm$ 2.0)
@@ V-A Training and Implementation Details
  • ER [12]: A rehearsal-based method that preserves a memory buffer containing samples from previous tasks and uses this buffer to facilitate the learning of new tasks. We impose a capacity limit of 1000 trajectories on the replay buffer.